OpenStudy (anonymous):

I need help really struggling understanding simplifying radical equations - square root of 252 x to 6th power y to 15th power

OpenStudy (anonymous):

\[\sqrt{256x^6y^{15}}\]is it yours?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what do you want to do,? simplifying it? first separate number and x, y into parts. tell me how can you take 256 out of the sqrt?

OpenStudy (anonymous):

yes simplify I think we are to find largest perfect square factor of 252

OpenStudy (anonymous):

and x and y varaible

OpenStudy (anonymous):

oh, sorry 252. ok, you pay attention at sqrt, sqrt mean exponent of 1/2 , right?

OpenStudy (anonymous):

take 252 divided by 2 how many did you get?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

you have 252 = 2*2*9*7 and then 2*2 = 2^2 9=3*3=3^2 so, 252 = 2^2*3*2*7, ok?